# SiamSeg: Self-Training with Contrastive Learning for Unsupervised Domain Adaptation Semantic Segmentation in Remote Sensing

### v2

@@ Abstract
Semantic segmentation of remote sensing (RS) images is a challenging yet crucial task. While deep learning, particularly supervised learning with large-scale labeled datasets, has significantly advanced this field, acquiring high-quality labeled data is expensive and time-consuming. Additionally, variations in ground sampling distance, imaging equipment, and geographic differences cause domain shifts between datasets, which limit model performance across domains. Unsupervised domain adaptation (UDA) offers a solution by enabling models to learn from unlabeled target domain data while training on labeled source domain data. Recent self-supervised learning approaches using pseudo-label generation have shown promise in addressing domain discrepancies. Combining source and target images with their true and pseudo labels has proven effective in reducing domain bias. However, the use of pseudo-labeling for RS image segmentation is underexplored. Existing methods often rely on high-confidence pixel points as pseudo-labels, reducing supervision in low-confidence areas. Noise in pseudo-labels further weakens the model’s ability to learn target domain semantics. While some methods assign confidence weights, noisy pseudo-labels remain an issue. To address these limitations, we propose integrating contrastive learning into UDA, enhancing the model’s capacity to capture semantic information by maximizing the similarity between augmented views of the same image. This provides additional supervision to improve performance in the target domain. Extensive experiments on key RS datasets (Potsdam, Vaihingen, LoveDA) demonstrate that our SimSeg method outperforms existing approaches, achieving state-of-the-art results. Visualization and quantitative analyses confirm its superior ability to learn from the target domain. The code is available at https://github.com/woldier/SiamSeg.
@@
Unsupervised Domain adaptation (UDA), Contrastive Learning, Remote Sensing (RS), Semantic Segmentation.
@@ I Introduction
Remote sensing techniques are widely employed in various visual tasks, including remote sensing scene classification [1] [2] [3] [4], target detection [5] [6] [7], and semantic segmentation [8] [9] [10] [11] [12]. Among these, the semantic segmentation task aims to accurately classify each pixel in remote sensing images for pixel-level target recognition. The extensive application of remote sensing image segmentation in urban planning, flood control, and environmental monitoring [13] has garnered increasing attention from researchers, prompting deeper exploration of the topic.
In recent years, deep learning-based semantic segmentation algorithms have made significant strides, leading to the emergence of many top-performing models, such as Fully Convolutional Neural Networks (FCNs) [14] [15] [16] and Transformer-based models [17] [18] [19]. However, the effectiveness of these methods heavily depends on the distributional similarity between the training and test data. When a domain shift occurs between different datasets, model performance significantly deteriorates. In practice, this domain shift problem is particularly pronounced due to the diversity of geographic regions, imaging conditions, and equipment used in remote sensing datasets, resulting in insufficient generalization capability of existing methods.
To address the domain bias problem and establish effective associations between source and target domains, cross-domain semantic segmentation of remote sensing images has emerged as a significant research direction. Unsupervised Domain Adaptation (UDA), a subset of transfer learning, aims to tackle the generalization challenge when the source domain has labeled data while the target domain contains only unlabeled data. Existing UDA methods can be broadly categorized into three groups: adversarial learning-based methods, generative model-based methods, and self-supervised learning (ST)-based methods [20].
Adversarial learning-based approaches assist segmentation networks in minimizing the differences between the distributions of feature maps in the source and target domains by introducing a discriminator [21] [22]. Specifically, the segmentation network first generates feature maps from input images in both domains and makes segmentation predictions. Meanwhile, the discriminator receives these feature maps as input and determines whether they originate from the source domain or the target domain. The segmentation network then aims to deceive the discriminator by adjusting the generated target domain feature maps to align closely with the source domain feature maps. Although adversarial learning methods have achieved some success in narrowing the distribution gap between source and target domains, they have significant drawbacks. First, the training process of adversarial learning is often unstable, particularly when dealing with complex feature spaces, making it prone to issues such as gradient vanishing or explosion. Second, as adversarial learning essentially involves a minimax game, the model may collapse during training; that is, the segmentation network may fail to generate diverse features and instead converge to a single solution, resulting in insufficient generalization ability. These challenges render adversarial learning methods difficult to apply in practical scenarios, particularly in cross-domain segmentation of remote sensing images.
In contrast to adversarial learning-based approaches, generative domain adaptation methods achieve visual similarity between source and target domain images at the pixel level using image-to-image translation or style migration techniques [23]. Specifically, these methods pass source domain images through a style translation network to generate images visually similar to the target domain. Subsequently, these style-migrated images are used to train a segmentation network alongside the labeled data from the source domain, improving performance on the target domain. Although generative methods minimize differences between source and target domains at the pixel level, they also have notable limitations. First, these methods typically require additional generative or style transformation networks, increasing model complexity and computational overhead. Second, the performance of the generative network directly impacts the final performance of the segmentation network. A style transformation network that fails to accurately capture the visual differences between the source and target domains or introduces noise and artifacts during the transformation process may hinder the segmentation network’s effectiveness on the target domain. Thus, this approach heavily relies on the conversion ability of the style transformation network, posing challenges for robustness and generalization in complex cross-domain scenarios.
Unlike adversarial learning and generative domain adaptation methods, self-training (ST) methods [13] do not rely on additional networks. The ST strategy facilitates cross-domain knowledge transfer by generating pseudo-labels [24] [25] [26] [27] [28]. The core idea is to follow the ”from easy to difficult” principle, prioritizing pixels with higher prediction confidence in the target domain as pseudo-label points. Labeled source domain data and pseudo-labeled target domain data are then input into the network for training, gradually improving model performance in the target domain. In remote sensing image segmentation, ST methods are relatively underexplored; this paper aims to apply pseudo-label-based ST methods to remote sensing semantic segmentation tasks. The advantages of ST methods include not needing additional generative networks and enhancing the target domain’s generalization ability through iterative optimization of pseudo-labels. Through this process, the pseudo-labels increasingly approximate the true distribution of the target domain, effectively improving segmentation network performance. However, challenges remain in the current ST approach. Most existing pseudo-label generation methods select only high-confidence pixel points for training, which, while ensuring the reliability of pseudo-labels to some extent, neglects low-confidence regions. This leads to insufficient supervised signals in these areas, preventing the network from fully learning the associated semantic information. Additionally, while some pseudo-labeling methods assign confidence weights to all pixels, they may introduce noise during pseudo-label generation, leading to misclassification of certain pixels. This noise weakens the supervised signals and can cause the network to learn inaccurate semantic information in the target domain. Consequently, although pseudo-labeling-based approaches partially address the domain adaptation challenge, they struggle with low-confidence regions and noisy labels, highlighting the urgent need for effective utilization of unlabeled target domain data.
@@
Figure 1: illustrates the main challenges in the task of semantic segmentation of remote sensing images across domains. These challenges include the problem of domain bias due to ground sampling distances, sensor differences, and variations in geographic landscapes, which affect the model’s ability to generalize across different datasets. From top to bottom, the first section shows the impact of different ground sampling distances from Potsdam IRRG to Vaihingen IRRG, the second section shows the remote sensing sensor variation from Potsdam RGB to Vaihingen IRRG, and the third section compares the geographic landscape differences between rural and urban areas in LoveDA. Details of the dataset will be discussed at IV-A. Dashed boxes in each case mark the differences in the same category across domains, highlighting the impact of these issues on semantic segmentation performance. Understanding these domain bias issues is crucial for improving the accuracy and robustness of semantic segmentation of remotely sensed images and points the way for subsequent research.
@@
More importantly, although many classical unsupervised domain adaptation (UDA) methods have been successfully applied to natural scenes, the domain shift problem in remote sensing images is more complex, stemming from factors such as ground sampling distances, sensor type discrepancies, and geographic landscape variations. As illustrated in Fig. 1, this complexity significantly degrades the performance of algorithms that work well in natural scenes when directly applied to remote sensing data. Specifically, self-supervised methods relying on pseudo-label generation show diminished richness in learned semantic features under complex domain shifts, resulting in lower confidence levels for generated pseudo-labels. The increased noise in these pseudo-labels raises the likelihood of pixel misclassification, further impairing the network’s ability to semantically characterize the target domain.
@@
The rise of contrast learning in computer vision demonstrates its powerful capability to capture semantic information in images without relying on labeled data, resulting in enhanced feature representation. Based on this observation, this paper proposes SiamSeg, which introduces contrast learning for the first time to the unsupervised domain adaptation task of semantic segmentation in remote sensing images. Leveraging the robust feature representation capability of contrast learning, SiamSeg effectively addresses the insufficient semantic information learning caused by the weak supervisory signals of pseudo-labels in the target domain, significantly enhancing segmentation network performance. To the best of our knowledge, SiamSeg is the first UDA approach to exploit contrast learning for facilitating context relations learning in the target domain.
@@
    Given the limited exploration of pseudo-labeled self-supervised learning in remote sensing UDA segmentation, this study focuses on the ST approach for UDA.
@@
    This paper presents the first application of contrast learning to an UDA task. Positive sample pairs are generated through different data augmentations of the same image, and the model is optimized using a Siamese network.
@@
  4. 4.
    Comprehensive experiments are conducted, demonstrating that the method achieves new state-of-the-art performance across various datasets
@@ II-A Unsupervised cross-domain adaptation for semantic segmentation
However, the sensitivity of the discriminator network to hyperparameters presents challenges in learning similar feature distributions across different domains. Additionally, adversarial learning may induce unstable training, further complicating the model’s learning process.
Another typical non-adversarial unsupervised domain adaptation paradigm is self-training, which has gained significant attention in cross-domain semantic segmentation in recent years. Zou et al. [26] first introduced a self-supervised learning method for unsupervised domain-adaptive semantic segmentation, wherein pseudo-labels are treated as discrete latent variables and computed by minimizing a uniform training objective. Tranheden et al., Zhou et al., Hoyer et al., and Chen et al. enhanced domain adaptation by generating trustworthy, consistent, and category-balanced pseudo-labels [27] [32] [33] [34].
However, there has been relatively limited exploration of self-supervised learning methods in the realm of semantic segmentation for remote sensing images. Therefore, this paper seeks to apply self-supervised learning to remote sensing image segmentation tasks. Despite the success of self-supervised learning methods in various applications, several challenges remain. For instance, current pseudo-label generation methods typically select only high-confidence pixels, leading to weakened supervised signals in low-confidence regions, which adversely affects the network’s ability to learn semantic information about the target domain. Additionally, self-supervised learning may introduce noise during the generation of pseudo-labels, resulting in some pixels being incorrectly assigned to erroneous pseudo-label categories. This further limits the model’s performance. In light of these challenges, integrating the advantages of contrastive learning may offer new perspectives for enhancing the application of self-supervised learning in remote sensing image segmentation.
@@ II-B Contrastive Learning
The core principle of contrastive learning is to generate pairs of images (view pairs or positive sample pairs) that share the same potential significance [35]. The optimization objective of contrastive learning is to encourage the model to learn similar embeddings for positive sample pairs while effectively distinguishing irrelevant sample pairs (negative sample pairs). This approach has gained prominence in unsupervised self-training representation learning [36] [37] [38]. The concept of simple and effective contrastive learning was further advanced through the introduction of the Siamese network [39] [40] [41] [42] [43] [44].
In practice, the performance of contrastive learning methods is significantly enhanced, partly due to the utilization of a large number of negative samples, which can be stored in a memory bank [36]. For instance, the MoCo method [41] maintains a queue of negative samples and employs a momentum encoder to improve the consistency of this queue. In contrast, the SimCLR method directly utilizes negative samples present in the current batch, though it typically requires a larger batch size to function effectively. The SimSiam method [44] achieves effective feature learning by simplifying the design. Unlike other contrastive learning methods, SimSiam does not rely on negative samples but instead builds “pairs of positive samples” for training. During training, the model processes different augmented versions of the same image and extracts features through a two-branch network. The two branches propagate forward separately, with a key component being the stop-gradient operation, which ensures that the parameters of one branch are not updated during backpropagation. This design facilitates the network’s learning capacity. SimSiam’s architecture allows it to perform well on small datasets, offering a new perspective on contrastive learning methods.
In the context of remote sensing (RS) images, the richer image features can lead to insufficient supervisory signals when using self-supervised learning with pseudo-labels. This limitation may hinder the effective learning of image features, adversely affecting the segmentation performance of the model. Therefore, enhancing the model’s representation learning ability through contrastive learning is essential, as it not only improves the accuracy of feature learning but also enhances the model’s performance in complex scenes.
@@ III-A Preliminary
In Unsupervised Domain Adaptation (UDA) for Remote Sensing (RS), we define two sets of images collected from different satellites or locations as distinct domains. To simplify the problem, we assume that the source domain and target domain images share the same pixel resolution, denoted as $H\times W\times 3$, where $H$ and $W$ represent the image height and width, respectively, and $3$ indicates the number of channels. Additionally, the two domains maintain consistency in the number of classes.
Let $x^{(i)}_{S}$ be the image and $y^{(i)}_{S}$ its corresponding label, with the source domain defined as $D_{S}=\left\{(x_{S}^{(i)},y_{S}^{(i)})\mid x_{S}^{(i)}\in\mathbb{R}^{H\times W\times 3},y_{S}^{(i)}\in\mathbb{R}^{H\times W\times C}\right\}_{i=1}^{N_{S}}$ , where $C$ is the number of classes. The target domain is defined as $D_{T}=\left\{x_{T}^{(i)}\mid x_{T}^{(i)}\in\mathbb{R}^{H\times W\times 3}\right\}_{i=1}^{N_{T}}$ , where only the images $x_{T}^{(i)}$ are accessible, while the labels $y_{T}^{(i)}$ remain unavailable. The subscripts $S$ and $T$ denote the source and target domains, respectively, and $N_{S}$ and $N_{T}$ indicate the number of samples in the source and target domains. It is important to note that during training, the labels $y_{T}^{(i)}$ from the target domain are inaccessible, being utilized only during the testing phase. The representation of the source domain label $y_{S}$ at the spatial position $(h,w)$ is denoted as a length $C$ one-hot encoding, represented as $y^{(i,h,w)}_{S}$, where $h\in[1,\dots,H]$ and $w\in[1,\dots,W]$.
The goal of UDA is to leverage the source domain images $x_{S}^{(i)}$ and their labels $y_{S}^{(i)}$ to train the model, enabling it to perform well on target domain images $x_{T}^{(i)}$ without access to the target domain labels $y_{T}^{(i)}$. If we solely rely on cross-entropy loss in the source domain for training the network $g_{\theta}$, it can be expressed as follows:
@@
In Equation (1), $g_{\theta}(x^{(i)}_{S})$ represents the predicted outcomes for each pixel in the source domain image $x^{(i)}_{S}$, providing normalized probabilities for each class at the spatial position $(h,w)$. However, due to the domain gap, relying solely on the source domain for training typically results in poor performance on the target domain, as the network struggles to generalize to target domain samples.
III-B Self-Training (ST) for UDA
To effectively address the domain gap, numerous methods have been proposed, broadly categorized into adversarial training [45] [29] [30] [31] and Self-Training (ST) [26] [27] [46]. Due to the instability of adversarial training, self-training has demonstrated superior performance across various tasks [27] [46]. Particularly in the field of remote sensing, research on ST methods is relatively scarce, which is why this study opts for a self-training approach.
To transfer knowledge from the source domain to the target domain, the ST method employs a teacher network $t_{\theta}$ to generate corresponding pseudo-labels for the target domain images (a detailed discussion on the teacher network will be presented in III-C). Mathematically, this is expressed in Equation (2):
@@ III-B Self-Training (ST) for UDA
where []delimited-[][$\cdot$][ ⋅ ] denotes the Iverson bracket. Here, $t_{\theta}(x_{T}^{(i)})$ indicates the class predictions for each pixel in the target domain image $x^{i}_{T}$. It is important to note that gradients are not backpropagated through the teacher network. Since we cannot ascertain the correctness of the generated pseudo-labels, it is necessary to evaluate the quality or confidence of the predictions for the pixels in the pseudo-labels. Only those pixels with class confidence exceeding a threshold $\tau$ will be used for training. The mathematical formulation for assessing the quality or confidence of the pixel at position $(h,w)$ is given by equation (3):
@@
The meaning of []delimited-[][$\cdot$][ ⋅ ] is consistent with that in Equation (2). We choose $t_{\theta}$ to generate pseudo-labels because it is more stable than the student network $g_{\theta}$. Higher thresholds generally correspond to higher-quality pseudo-labels; however, this may result in fewer pseudo-label pixels being utilized for training during the early phases. As the training iterations increase, confidence gradually rises, leading to an increase in the number of pseudo-labels employed.
In addition to training on labeled data in the source domain using Equation (1), the pseudo-labels $p_{T}^{(i)}$ and their corresponding quality estimates $q_{T}^{(i)}$ from the target domain will also be incorporated into the training of the student network $g_{\theta}$. The loss for training in the target domain can be mathematically represented as Equation (4):
@@
Here, $g_{\theta}(x^{(i)}_{T})$ denotes the predicted outcomes for each pixel in the target domain image $x^{(i)}_{T}$, providing normalized probability predictions. Through this approach, the model can effectively utilize pseudo-label information to enhance its performance in the target domain.
III-C Pseudo-Label Generation
The generation of pseudo-labels $p_{T}^{(i)}$ can be conducted through online [47] [27] or offline [25] [48] [26] methods. In online self-training, the teacher network $t_{\theta}$ updates in tandem with $g_{\theta}$ during the training process. This procedure is relatively straightforward; however, it is crucial to note that the structures of the teacher network $t_{\theta}$ and the student network $g_{\theta}$ are identical. Moreover, due to the use of the stop-gradient operation, the teacher network does not backpropagate gradients during training. Typically, to mitigate the parameter perturbations caused by the domain gap during the training process, the weights of the teacher network are updated after each training iteration using the exponentially moving average (EMA) method [49], thereby enhancing the stability of pseudo-label generation. This can be mathematically expressed as Equation (5):
@@ III-C Pseudo-Label Generation
where $S_{t}(\cdot)$ denotes the weights of the model at training step $t$, and the hyperparameter $\alpha\in[0,1]$ indicates the importance of the current state $S(t_{\theta})$. This weight update strategy not only enhances the accuracy of pseudo-label generation but also alleviates the adverse effects of domain differences, ultimately improving the model’s generalization ability in the target domain.
While pseudo-labeling methods have made significant progress in addressing domain adaptation, they still face challenges. In particular, noisy pseudo-labels can arise from low-confidence regions, leading to pixel misclassification and weakening the supervised signals needed for effective learning in the target domain. As domain shifts in remote sensing images become more complex—due to factors like sensor discrepancies and variations in geographic landscapes—the accuracy of pseudo-labels tends to degrade. This increased noise introduces additional difficulty in training the model to semantically characterize the target domain, limiting the overall effectiveness of these methods.
III-D Contrastive Learning
To overcome the limitations of pseudo-labeling, we introduce a contrastive learning. The contrastive learning module generates two distinct views $x_{T}^{(i)}$ through two random augmentations. These views are processed by an encoder network $E$ , consisting of a backbone $f$ (e.g., MIT [18]) and a Multi-Layer Perceptron (MLP) projection head [42]. The weights of the encoder network $E$ remain the same while processing both views. We denote the MLP prediction head as $h$; the output of one view through the encoder network $E$ is transformed to match the representation of the other view. The outputs of the MLP prediction head and the MLP projection head can be expressed as $p_{1}^{(i)}=h(E(x_{T_{1}}^{(i)}))$ and $z_{1}^{(i)}=E(x_{T_{2}}^{(i)})$. Our objective is to optimize the negative cosine similarity between these two vectors, mathematically represented as Equation (6):

### v3

@@ Abstract
Semantic segmentation of remote sensing (RS) images is a challenging yet essential task with broad applications. While deep learning, particularly supervised learning with large-scale labeled datasets, has significantly advanced this field, the acquisition of high-quality labeled data remains costly and time-intensive. Unsupervised domain adaptation (UDA) provides a promising alternative by enabling models to learn from unlabeled target domain data while leveraging labeled source domain data. Recent self-training (ST) approaches employing pseudo-label generation have shown potential in mitigating domain discrepancies. However, the application of ST to RS image segmentation remains underexplored. Factors such as variations in ground sampling distance, imaging equipment, and geographic diversity exacerbate domain shifts, limiting model performance across domains. In that case, existing ST methods, due to significant domain shifts in cross-domain RS images, often underperform. To address these challenges, we propose integrating contrastive learning into UDA, enhancing the model’s ability to capture semantic information in the target domain by maximizing the similarity between augmented views of the same image. This additional supervision improves the model’s representational capacity and segmentation performance in the target domain. Extensive experiments conducted on RS datasets, including Potsdam, Vaihingen, and LoveDA, demonstrate that our method, SimSeg, outperforms existing approaches, achieving state-of-the-art results. Visualization and quantitative analyses further validate SimSeg’s superior ability to learn from the target domain. The code is publicly available at https://github.com/woldier/SiamSeg.
@@
Unsupervised Domain Adaptation, Contrastive Learning, Remote Sensing, Semantic Segmentation.
@@ I Introduction
Figure 1: The main challenges in the task of cross-domain semantic segmentation of remote sensing images. These challenges include the problem of domain bias due to ground sampling distances, sensor differences, and variations in geographic landscapes, which affect the model’s ability to generalize across different datasets. Understanding these domain shift issues is crucial for improving the accuracy and robustness of semantic segmentation of RS images.
Remote sensing techniques are widely employed in various visual tasks, including RS images classification [1, 2, 3, 4], object detection [5, 6, 7], and semantic segmentation [8, 9, 10, 11, 12, 13]. Among these, the semantic segmentation task aims to accurately classify each pixel in remote sensing images for pixel-level target recognition. The extensive application of remote sensing image segmentation in urban planning, flood control, and environmental monitoring [14] has garnered increasing attention from researchers, prompting deeper exploration of the topic.
In recent years, deep learning-based semantic segmentation methods have made significant strides, leading to the emergence of many top-performing models, such as Fully Convolutional Neural Networks [15, 16, 17] and Transformer-Based Models [18, 19, 20]. However, the effectiveness of these methods heavily depends on the distributional similarity between the training and test data. When a domain shift occurs between different datasets, model performance significantly deteriorates. In practice, this domain shift problem is particularly pronounced due to the diversity of geographic regions, imaging conditions, and equipment used in remote sensing datasets, resulting in insufficient generalization capability of existing methods.
To address the domain shift problem and establish effective associations between source and target domains, cross-domain semantic segmentation of remote sensing images has emerged as a significant research direction. Unsupervised domain adaptation (UDA), a subset of transfer learning, aims to tackle the generalization challenge when the source domain has labeled data while the target domain contains only unlabeled data. Existing UDA methods can be broadly categorized into two groups: adversarial learning based methods and self-training (ST) based methods [21]. Adversarial learning based approaches assist segmentation networks in minimizing the differences between the distributions of feature maps in the source and target domains by introducing a discriminator [22, 23, 24, 25]. Unlike adversarial learning and generative domain adaptation methods, self-training (ST) methods [14] do not rely on additional discriminators. The ST strategy facilitates cross-domain knowledge transfer by generating pseudo-label [26, 27, 28, 29, 30].
@@
Although many classical UDA methods have been successfully applied to natural scenes, the domain shift problem in remote sensing images is more complex, as illustrated in Fig. 1, stemming from factors such as ground sampling distances, sensor type discrepancies, and geographic landscape variations. This results in a larger domain gap for cross-domain RS images and significantly degrades the performance of methods that work well in natural scenes when directly applied to remote sensing data.
@@
Directly applying the ST method to cross-domain RS image semantic segmentation does not capture the feature information of the target domain image well, which leads to the performance degradation of the ST method in the target domain. The rise of contrast learning in computer vision demonstrates its powerful capability to capture semantic information in images without relying on labeled data, resulting in enhanced feature representation. This addresses the issue of a large domain gap, which prevents the application of ST methods to learn the target domain effectively through pseudo-label. Based on this observation, this paper proposes SiamSeg, which introduces contrast learning to the unsupervised domain adaptation task of semantic segmentation in remote sensing images. Leveraging the robust feature representation capability of contrast learning, SiamSeg effectively addresses the insufficient semantic information learning caused by the weak supervisory signals of pseudo-label in the target domain, significantly enhancing segmentation network performance.
@@
    Given the limited exploration of ST in RS UDA segmentation, this study focuses on the ST approach for UDA.
@@
    Due to the large domain gap of cross-domain RS image, the existing ST methods cannot learn the features of the target domain well. Therefore This paper presents the first application of contrast learning to an UDA task.
@@ II-A Unsupervised cross-domain adaptation for semantic segmentation
Another typical non-adversarial unsupervised domain adaptation paradigm is self-training (ST), which has gained significant attention in cross-domain semantic segmentation in recent years. Zou et al. [28] first introduced a ST method for unsupervised domain-adaptive semantic segmentation. Tranheden et al. [29], Zhou et al. [34] , Hoyer et al. [35], and Chen et al [36]. enhanced domain adaptation by generating trustworthy, consistent, and category-balanced pseudo-label.
@@ II-B Contrastive Learning
However, as shown in Fig. 1, compared with natural images in cross-domain RS images the domain gap is larger. Directly applying the ST method to cross-domain RS image semantic segmentation does not capture the feature information of the target domain image well, which leads to the performance degradation of the ST method in the target domain. Therefore, this paper tries to use contrast learning to make up for this defect.
The core principle of contrastive learning is to generate pairs of images (view pairs or positive sample pairs) that share the same potential significance [37]. The optimization objective of contrastive learning is to encourage the model to learn similar embeddings for positive sample pairs while effectively distinguishing irrelevant sample pairs (negative sample pairs). This approach has gained prominence in unsupervised self-training representation learning [38, 39, 40]. The concept of simple and effective contrastive learning was further advanced through the introduction of the Siamese network [41, 42, 43, 44, 45, 46].
In practice, the performance of contrastive learning methods is significantly enhanced, partly due to the utilization of a large number of negative samples, which can be stored in a memory bank [38]. For instance, the MoCo method [43] maintains a queue of negative samples and employs a momentum encoder to improve the consistency of this queue. In contrast, the SimCLR method directly utilizes negative samples present in the current batch, though it typically requires a larger batch size to function effectively. The SimSiam method [46] achieves effective feature learning by simplifying the design. Unlike other contrastive learning methods, SimSiam does not rely on negative samples but instead builds “pairs of positive samples” for training. In the context of remote sensing (RS) images, the richer image features and larger domain gap can lead to insufficient feature learning when using ST method. This limitation may hinder the effective learning of image features, adversely affecting the segmentation performance of the model. Therefore, enhancing the model’s representation learning ability through contrastive learning is essential, as it not only improves the accuracy of feature learning but also enhances the model’s performance in complex scenes.
@@ III-A Preliminary
In Unsupervised domain adaptation (UDA) for remote sensing (RS), we define two sets of images collected from different satellites or locations as distinct domains. To simplify the problem, we assume that the source domain and target domain images share the same pixel resolution, denoted as $H\times W\times 3$. Additionally, the two domains maintain consistency in the number of classes.
Let $x^{(i)}_{S}$ be the image and $y^{(i)}_{S}$ its corresponding label, with the source domain defined as $D_{S}=\left\{(x_{S}^{(i)},y_{S}^{(i)})\mid x_{S}^{(i)}\in\mathbb{R}^{H\times W\times 3},y_{S}^{(i)}\in\mathbb{R}^{H\times W\times C}\right\}_{i=1}^{N_{S}}$ , where $C$ is the number of classes. The target domain is defined as $D_{T}=\left\{x_{T}^{(i)}\mid x_{T}^{(i)}\in\mathbb{R}^{H\times W\times 3}\right\}_{i=1}^{N_{T}}$ , where only the images $x_{T}^{(i)}$ are accessible, while the labels $y_{T}^{(i)}$ remain unavailable. The subscripts $S$ and $T$ denote the source and target domains, respectively, and $N_{S}$ and $N_{T}$ indicate the number of samples in the source and target domains. The representation of the source domain label $y_{S}$ at the spatial position $(h,w)$ is denoted as a length $C$ one-hot encoding, represented as $y^{(i,h,w)}_{S}$, where $h\in[1,\dots,H]$ and $w\in[1,\dots,W]$.
If we solely rely on cross-entropy loss in the source domain for training the network $g_{\theta}$, it can be expressed as follows:
@@
In Equation (1), $g_{\theta}(x^{(i)}_{S})$ represents the predicted outcomes for each pixel in the source domain image $x^{(i)}_{S}$. However, due to the domain gap, relying solely on the source domain for training typically results in poor performance on the target domain, as the network struggles to generalize to target domain samples.
III-B Self-Training for UDA
To transfer knowledge from the source domain to the target domain, the ST method employs a teacher network $t_{\theta}$ to generate corresponding pseudo-label for the target domain images. Mathematically, this is expressed in Equation (2):
@@ III-B Self-Training for UDA
where []delimited-[][$\cdot$][ ⋅ ] denotes the Iverson bracket. Here, $t_{\theta}(x_{T}^{(i)})$ indicates the class predictions for each pixel in the target domain image $x^{(i)}_{T}$. It is important to note that gradients are not backpropagated through the teacher network. Since we cannot ascertain the correctness of the generated pseudo-label, it is necessary to evaluate the quality or confidence of the predictions for the pixels in the pseudo-label. Only those pixels with class confidence exceeding a threshold $\tau$ will be used for training. The mathematical formulation for assessing the quality or confidence of the pixel at position $(h,w)$ is given by equation (3):
@@
In addition to training on labeled data in the source domain using Equation (1), the pseudo-label $p_{T}^{(i)}$ and their corresponding quality estimates $q_{T}^{(i)}$ from the target domain will also be incorporated into the training of the student network $g_{\theta}$. The loss for training in the target domain can be mathematically represented as Equation (4):
@@
During the training process, the weights of the teacher network are updated after each training iteration using the exponentially moving average (EMA) method [47], thereby enhancing the stability of pseudo-label generation. This can be mathematically expressed as Equation (5):
@@
where $S_{t}(\cdot)$ denotes the weights of the model at training step $t$, and the hyperparameter $\alpha\in[0,1]$ indicates the importance of the current state $S(t_{\theta})$.
III-C Contrastive Learning
In the context of RS images, the richness of image features can exacerbate domain gaps and lead to insufficient feature learning when applying the ST method. These limitations may hinder effective feature learning, ultimately degrading the model’s segmentation performance. To overcome the limitations of ST methods in RS, we introduce a contrastive learning to improve methods performance in the target domain. The contrastive learning module generates two distinct views $x_{T}^{(i)}$ through two random augmentations. These views are processed by an encoder network $E$ , consisting of a backbone $f$ (e.g., MIT [19]) and a Multi-Layer Perceptron (MLP) projection head [44]. The weights of the encoder network $E$ remain the same while processing both views. We denote the MLP prediction head as $h$; the output of one view through the encoder network $E$ is transformed to match the representation of the other view. The outputs of the MLP prediction head and the MLP projection head can be expressed as $p_{1}^{(i)}=h(E(x_{T_{1}}^{(i)}))$ and $z_{1}^{(i)}=E(x_{T_{2}}^{(i)})$. Our objective is to optimize the negative cosine similarity between these two vectors, mathematically represented as Equation (6):